A population is given by 5 8 10 15 16 20 draw all possible samples of size 2 without replacement. Show that E(x ̅ )=μ


1
Expert's answer
2021-11-22T13:14:24-0500

N=6

n=2

"\mu = \frac{5+ 8+ 10+ 15+ 16+ 20}{6} = 12.33"

Number of samples "= \frac{6!}{2!(6-2)!} = \frac{5 \times 6}{2} = 15"


"E(\bar{x}) = \frac{6.5+7.5+...+17.5+18}{15}=12.33"